Two pendulums of time periods 4s and 9 s respectively start oscillating simultaneously from two opposite extreme positions. After how much time they will be in the same phase?

Grade:12

1 Answers

Arbind
6 years ago
Let n is number of oscillation of smaller pendulum then, nx4=(n-1)x9,n=9/5Time = (9/5)x4= 7.2 seconds